# FIGRET: Fine-Grained Robustness-Enhanced Traffic Engineering

### v1

@@ Abstract.
Traffic Engineering (TE) is critical for improving network performance and reliability. A key challenge in TE is the management of sudden traffic bursts. Existing TE schemes often struggle to accurately determine the extent of focus required for these surges, thereby facing difficulties in achieving a balance between performance under normal and peak traffic conditions. To address this issue, we introduce FIGRET, a Fine-Grained Robustness-Enhanced TE Scheme. FIGRET offers a novel approach to TE by providing varying levels of robustness enhancements, customized according to the distinct traffic characteristics of various source-destination pairs. By leveraging a sophisticated loss function and advanced deep learning techniques, FIGRET is capable of generating high-quality TE solutions efficiently. Our evaluations of real-world production networks, including Wide Area Networks and data centers, demonstrate that FIGRET significantly outperforms existing TE schemes. Compared to the TE scheme currently deployed in Google’s Jupiter network, FIGRET achieves a 9%-34% reduction in average Maximum Link Utilization and improves solution speed by 35×35$\times$35 ×-1800×1800$\times$1800 ×. Against DOTE, a state-of-the-art deep learning-based TE method, FIGRET substantially lowers the occurrence of significant congestion events triggered by traffic bursts by 41%-53.9% in topologies characterized by high traffic dynamics.
copyright: noneconference: ACM Conference; 2024; Sydney, Australiadoi: isbn: price: footnotetext: *Shizhen Zhao is the corresponding author.
@@
With the exponential growth in network traffic, both data centers (Poutievski et al., 2022; Alizadeh et al., 2014; Benson et al., 2011; Al-Fares et al., 2010; Teh et al., 2020; Cao et al., 2021; Chen et al., 2018) and Wide Area Networks (WANs) (Jiang et al., 2009; Kandula et al., 2005; Kandula et al., 2014; Kumar et al., 2018b; Elwalid et al., 2001; Fortz and Thorup, 2000; Zhong et al., 2021; Wang et al., 2006; Roughan et al., 2003; Applegate and Cohen, 2003; Suchara et al., 2011; Xu et al., 2023) are increasingly dependent on Traffic Engineering (TE) for optimizing network flow. TE, typically enabled by a centralized controller in Software-Defined Networking (SDN) (Jain et al., 2013; Poutievski et al., 2022; Akyildiz et al., 2014; Agarwal et al., 2013; Zaicu et al., 2021; Bogle et al., 2019; Hong et al., 2013; Kumar et al., 2015; Liu et al., 2014), periodically solves optimization problems to efficiently route traffic across network paths, translating these solutions into router configurations.
@@
A major challenge in TE is managing sudden traffic bursts. Considering the delays introduced by the central controller in collecting traffic demands, computing new TE solutions, and updating forwarding rules, TE systems should predefine network configurations based on historical data prior to the arrival of actual traffic flows. Yet, the inherently dynamic and unpredictable nature of real network traffic poses significant forecasting difficulties (Teixeira et al., 2005; Xu et al., 2005). Inadequate preparation for these bursts can result in severe network congestion, leading to lengthy delays, high packet loss rates, and diminished network throughput (Wang et al., 2006). Thus, enhancing the robustness to respond to unexpected traffic bursts is essential.
Existing TE schemes demonstrate limitations in handling traffic bursts, struggling to determine the appropriate level of intervention for such scenarios. One TE approach to managing traffic bursts is oblivious routing (Applegate and Cohen, 2003), which involves optimizing TE solutions for the worst-case scenario across all traffic demands. While this equips the system to handle bursts, it tends to result in suboptimal performance for typical traffic, which constitutes the majority of traffic over time. Consequently, alternative TE methods (Wang et al., 2006; Poutievski et al., 2022; Teh et al., 2020, 2022) do not focus on optimizing for all traffic demands. Instead, they design TE solutions based on historical data while imposing certain constraints to enhance robustness. For example, Google’s hedging mechanism (Poutievski et al., 2022) introduces a metric to evaluate the impact of burst traffic on each path, requiring that in solutions, the metric for each path must not surpass a predetermined upper limit. Since optimization occurs based on historical traffic matrices, the impact on average performance is less significant compared to oblivious routing. However, depending solely on a single constraint for optimization often fails to achieve a satisfactory balance in TE performance across regular and bursty traffic conditions.
@@
The limitations observed in previous methods can be attributed to their excessively broad approach to enhancing robustness. These methods adopt a perspective grounded in a traffic matrix that encompasses flows from all source-destination pairs across the entire network. In doing so, they treat the traffic of each pair indiscriminately and uniformly. However, in practice, network traffic exhibits significant variations among different source-destination pairs. While some pairs may encounter traffic bursts, others may maintain remarkable stability. For consistently stable traffic, prioritizing robustness becomes unnecessary and may even result in performance compromises during regular traffic conditions.
@@
Building upon this observation, we design FIGRET (Fine-Grained Robustness-Enhanced Traffic Engineering). FIGRET’s key insight lies in a fine-grained customized robustness enhancement strategy, tailored for individual source-destination pairs based on their traffic characteristics. It applies relaxed robustness requirements to source-destination pairs with stable traffic and enforces stricter requirements for those prone to bursts. In the initial phase, we employ a metric called path sensitivity (Teh et al., 2020) to measure the impact of burst traffic on network paths, thus enabling robustness enhancement through the application of path sensitivity constraints. To achieve fine-grained robustness and ensure a satisfactory balance in TE performance across both normal and bursty traffic conditions, we customize the path sensitivity constraints according to the network topology and the traffic characteristics of the source-destination pairs they connect, while also considering this path sensitivity constraint when optimizing based on historical data. To accomplish this, FIGRET leverages a deep neural network equipped with a well-designed loss function, which includes two components: one for evaluating performance under normal traffic conditions and another for assessing robustness. Moreover, we employ an efficient network architecture for FIGRET.
@@
We conducted a comprehensive evaluation of FIGRET. This evaluation utilized publicly available WAN datasets (Uhlig et al., 2006), as well as Data Center PoD-level and ToR-level topologies and traffic data (Alizadeh et al., 2013; Roy et al., 2015). The data encompassed topologies ranging from dozens to hundreds of nodes, with the corresponding traffic data exhibiting various characteristics, including stable, composite, and bursty profiles. Through our evaluation, we found that FIGRET consistently delivers high-quality TE solutions across a variety of topologies. Compared to the TE system currently in production by Google’s Jupiter (Poutievski et al., 2022), FIGRET achieved an average Max Link Utilization (MLU) reduction of 9%-34% across different topologies and improved solution speed by 35×35$\times$35 ×-1800×1800$\times$1800 ×. In comparison with the state-of-the-art Deep Learning-based TE system DOTE (Perry et al., 2023), specifically designed for MLU optimization, FIGRET achieved notable improvements in two topologies with bursty traffic data. It reduced the average Max Link MLU by 4.5% and 5.3% while simultaneously decreasing the number of significant congestion incidents caused by traffic bursts by 41% and 53.9%. Meanwhile, in topologies with stable traffic data, FIGRET performed at least as well as DOTE, despite the additional consideration of robustness. Additionally, we provided interpretability for the workings of FIGRET.
@@
To motivate our fine-grained robustness enhancement method, we first demonstrate that different source-destination pairs within a network exhibit distinct traffic characteristics (§2.1). Subsequently, we illustrate the necessity for robustness in TE correlates with traffic characteristics. By tailoring robustness requirements according to the traffic characteristics, TE becomes more effective than treating all source-destination pairs as equal (§2.2).
@@ 2.1. Diversity in traffic characteristics
To demonstrate the diverse traffic characteristics under different source-destination pairs (SD pairs), we provide an analysis of traffic characteristics across SD pairs in various production networks. This includes the WAN network GEANT (Uhlig et al., 2006), as well as the Facebook data center network (Roy et al., 2015) at both the PoD-level and ToR-level. Details on the public data used are provided in §5. For each SD pair, we calculate the variance as a measure of the dynamic changes in traffic. The results of our analysis are displayed in Figure 1, which shows the variance of traffic demand for different SD pairs across the three types of network traffic. The greater the variance for a SD pair, the more unstable that SD pair is.
(a) PoD-level, Facebook DB cluster
(b) WAN GEANT
(c) ToR-level, Facebook DB cluster
Figure 1. The variance of traffic demand by source and destination (The value of variance has been normalized). The larger the variance, the more likely it is that the traffic demand for that source-destination pair will experience a burst. Regardless of which network in 1(a), 1(b), or 1(c), the traffic demands of different source-destination pairs exhibit different traffic characteristics.
The results in Figure 1 indicate that, regardless of the topology type, there are distinct traffic characteristics for different SD pairs. It is inappropriate to consider all traffic within these SD pairs as uniform, as this approach inevitably leads to a trade-off in TE. If TE treats all SD pairs as non-bursty traffic, it sacrifices the ability of TE to handle bursty traffic effectively. On the other hand, treating all SD pairs as bursty traffic results in suboptimal performance under non-bursty conditions. So, we seek a method that can differently consider SD pairs with varying traffic characteristics.
@@ 2.2. Dig deep into the trade-off
TE scheme 1 considers all traffic as non-bursty and only optimizes for the typical situation. To minimize congestion in non-bursty scenarios, TE scheme 1 directs all traffic along the shortest paths. In the typical situation, its Max Link Utilization (MLU) is $\max\{\frac{1}{2},\frac{1}{2},\frac{1}{2}\}=0.5$. When any burst situation occurs, the MLU is $\max\{\frac{4}{2},\frac{1}{2},\frac{1}{2}\}=2$.
TE scheme 2 considers all traffic as bursty traffic and aims to enhance robustness in response to the three burst situations by splitting traffic across different paths. When dealing with the typical situation, the MLU of TE scheme 2 is $\max\{\frac{1\times 0.5+1\times 0.5+1\times 0.5}{2},\frac{1\times 0.5+1\times 0.5+1\times 0.5}{2},\frac{1\times 0.5+1\times 0.5+1\times 0.5}{2}\}=0.75$, while dealing with the three burst situations, its MLU is $\max\{\frac{4\times 0.5+1\times 0.5+1\times 0.5}{2},\frac{4\times 0.5+1\times 0.5+1\times 0.5}{2},\frac{4\times 0.5+1\times 0.5+1\times 0.5}{2}\}=1.5$. That is, TE scheme 2 exhibits better robustness in handling traffic bursts compared to TE scheme 1, but with a decrease in typical performance.
Overall, this one-size-fits-all approach either leads to a network lacking the capacity to handle bursts or enhances the capability to manage bursts but at the cost of compromised performance in non-burst situations. This trade-off presents a dilemma for TE to address.
@@
Based on insights from Figure 1 and Figure 2, it is evident that enhancing robustness to manage unexpected traffic bursts is necessary. However, uniformly increasing the robustness across all paths may lead to a compromise in average performance. A more nuanced strategy would be to adjust the robustness of paths based on the traffic characteristics of the traffic demands they serve. This strategy, which refines the robustness enhancement at a finer granularity compared to previous TE methods, enables a better balance between typical performance and robustness in TE.
(a) Network traffic pattern.
(b) Topology
(c) TE scheme 1
(d) TE scheme 2
(e) TE scheme 3
Figure 2. A simple example comparing three TE schemes. TE scheme 1 has the optimal MLU in the typical situation. When traffic burst situations 1/2/3 are all possible, TE scheme 2 exhibits the best resilience against bursts. However, if only traffic burst situation 3 is likely to occur, TE scheme 3 outperforms TE scheme 2.
@@ 3. TE model
  • Network paths. Each source vertex $s$ communicates with each destination vertex $d$ via a set of network paths $P_{sd}$. We can also say that $P_{sd}$ serves the SD pair.
  • Path capacity. The capacity of a path $p$ is denoted by $C_{p}$. It is the minimum capacity across the edges comprising the path (Poutievski et al., 2022).
  • TE configuration. A TE configuration $\mathcal{R}$ specifies for each source vertex $s$ and destination vertex $d$ how the $D_{sd}$ traffic from $s$ to $d$ is split across the paths in $P_{sd}$. A TE configuration specifies for each path $p\in P_{sd}$ a split ratio $r_{p}$, where $r_{p}$ is the fraction of the traffic demand from $s$ to $d$ forwarded along path $p$ ($\sum_{p\in P_{sd}}r_{p}=1$).
@@
    Given a demand matrix $D$ and TE configuration $\mathcal{R}$, the traffic flow traversing edge $e$ is $f_{e}=\sum_{s,d\in V,p\in P_{sd},e\in p}D_{sd}\cdot r_{p}$, then the MLU induced by $D$ and $\mathcal{R}$ is $\text{MLU}=\max_{e\in E}\frac{f_{e}}{c(e)}$, which we denote as $M(\mathcal{R},D)$.
TE under traffic uncertainty. At epoch $t$, the decision maker in TE must output a network configuration $\mathcal{R}_{t}(D_{1},\dots,D_{t-1})$ based on historical data $\{D_{1},\dots,D_{t-1}\}$ before the arrival of $D_{t}$. The goal is to ensure that this configuration consistently yields low link over-utilization for the upcoming traffic $D_{t}$.
@@
When configuring $\mathcal{R}_{t}$ based on historical data $\{D_{1},\dots,D_{t-1}\}$ before the arrival of $D_{t}$, it is critical to consider both anticipated and unanticipated traffic. Firstly, an anticipated traffic demand $D_{t}^{\text{expect}}$ is deduced from historical traffic data. This expected demand may be explicitly obtained through predictive modeling or implicitly represent the insights gleaned from past data trends. Concurrently, acknowledging the inevitable discrepancies between anticipated and actual traffic due to mis-prediction, it is crucial to incorporate the impact of such mismatches when calculating the TE configuration.

### v3

@@ Abstract.
Traffic Engineering (TE) is critical for improving network performance and reliability. A key challenge in TE is the management of sudden traffic bursts. Existing TE schemes either do not handle traffic bursts or uniformly guard against traffic bursts, thereby facing difficulties in achieving a balance between normal-case performance and burst-case performance. To address this issue, we introduce FIGRET, a Fine-Grained Robustness-Enhanced TE scheme. FIGRET offers a novel approach to TE by providing varying levels of robustness enhancements, customized according to the distinct traffic characteristics of various source-destination pairs. By leveraging a burst-aware loss function and deep learning techniques, FIGRET is capable of generating high-quality TE solutions efficiently. Our evaluations of real-world production networks, including Wide Area Networks and data centers, demonstrate that FIGRET significantly outperforms existing TE schemes. Compared to the TE scheme currently deployed in Google’s Jupiter data center networks, FIGRET achieves a 9%-34% reduction in average Maximum Link Utilization and improves solution speed by 35×35$\times$35 ×-1800×1800$\times$1800 ×. Against DOTE, a state-of-the-art deep learning-based TE method, FIGRET substantially lowers the occurrence of significant congestion events triggered by traffic bursts by 41%-53.9% in topologies with high traffic dynamics.
Traffic engineering, Wide-Area Networks, Datacenter networks, Machine learning
journalyear: 2024copyright: rightsretainedconference: ACM SIGCOMM 2024 Conference; August 4–8, 2024; Sydney, NSW, Australiabooktitle: ACM SIGCOMM 2024 Conference (ACM SIGCOMM ’24), August 4–8, 2024, Sydney, NSW, Australiadoi: 10.1145/3651890.3672258isbn: 979-8-4007-0614-1/24/08ccs: Networks Traffic engineering algorithmsccs: Computing methodologies Machine learning
@@
With the exponential growth in network traffic, both data center networks (Poutievski et al., 2022; Alizadeh et al., 2014; Benson et al., 2011; Al-Fares et al., 2010; Teh et al., 2020; Cao et al., 2021; Chen et al., 2018) and Wide Area Networks (WANs) (Jiang et al., 2009; Kandula et al., 2005, 2014; Kumar et al., 2018; Elwalid et al., 2001; Fortz and Thorup, 2000; Zhong et al., 2021; Wang et al., 2006; Roughan et al., 2003; Applegate and Cohen, 2003; Suchara et al., 2011; Xu et al., 2023) are increasingly dependent on Traffic Engineering (TE) to optimize network performance. TE, typically enabled by a centralized controller in Software-Defined Networking (SDN) (Jain et al., 2013; Poutievski et al., 2022; Akyildiz et al., 2014; Agarwal et al., 2013; Zaicu et al., 2021; Bogle et al., 2019; Hong et al., 2013; Kumar et al., 2015; Liu et al., 2014), periodically solves optimization problems to efficiently route traffic across network paths, and then translates these solutions into router configurations.
@@
A major challenge in TE is managing sudden traffic bursts. Considering the latency introduced by the central controller in collecting traffic demands, computing new TE solutions, and updating forwarding rules, A TE system should pre-compute network configurations based on historical data prior to the arrival of actual traffic flows. Yet, the inherently dynamic and unpredictable nature of real network traffic poses significant forecasting difficulties (Teixeira et al., 2005; Xu et al., 2005). Inadequate preparation for traffic bursts may result in severe network congestion, leading to prolonged delays, high packet loss rates, and diminished network throughput (Wang et al., 2006). Thus, enhancing the robustness against unexpected traffic bursts is essential.
Existing burst-aware TE schemes usually handle traffic bursts at the cost of compromised normal-case network performance. Oblivious routing  (Applegate and Cohen, 2003) optimizes TE solutions for the worst-case scenario across all traffic demands. While this approach offers the highest level of robustness against traffic bursts, it tends to result in highly suboptimal performance for non-burst traffic patterns, which constitute the majority of traffic over time. As an improvement, Cope (Wang et al., 2006) focuses on optimizing the predicted traffic demands while providing a worst-case performance guarantee. However, offering such a worst-case guarantee may still be an overkill because some traffic patterns may never occur. In addition, the computational complexity of COPE is also much higher. Consequently, a new class of TE methods (Wang et al., 2006; Poutievski et al., 2022; Teh et al., 2020, 2022) are proposed. These TE methods do not offer guarantee in the entire traffic pattern space, but instead enhance robustness by directly limiting the routing weights of different paths. Specifically, COUDER (Teh et al., 2022) introduces the path sensitivity metric to evaluate the impact of burst traffic on each path, and enhances robustness by minimizing the maximum sensitivity across all paths. Similarly, Google’s optical data centers adopt a hedging mechanism in TE(Poutievski et al., 2022), which enhances robustness by confining the path sensitivity under a predetermined upper limit. However, these methods may still impact TE performance under non-burst traffic scenarios because they force traffic to spread across multiple (potentially longer) paths rather than taking the best path, even if the traffic between certain source-destination pairs is stable.
@@
The limitations observed in the previous methods can be attributed to their uniform treatment against traffic bursts. In practice, network traffic exhibits different characteristics among different source-destination pairs. While some pairs may frequently encounter traffic bursts, others could remain remarkably stable. For consistently stable traffic, prioritizing robustness becomes unnecessary and may even hurt performance.
@@
Building upon this observation, we design FIGRET (Fine-Grained Robustness-Enhanced Traffic Engineering). FIGRET’s key insight lies in a fine-grained customized robustness enhancement strategy, tailored for every source-destination pair based on their traffic characteristics. It applies relaxed robustness requirements to source-destination pairs with stable traffic and enforces stricter requirements for those prone to bursts. Similar to (Teh et al., 2020), FIGRET also uses the path sensitivity metric to enhance robustness against traffic bursts. On top of this, FIGRET customizes the path sensitivity constraints according to the network topology and the traffic characteristics of different source-destination pairs. This strategy allows FIGRET to achieve fine-grained robustness enhancement and ensure a satisfactory balance in TE performance between both normal and burst traffic scenarios.
Having proposed the FIGRET formulation, the next challenge is to efficiently compute a TE solution. At first glance, FIGRET’s formulation can be directly solved using linear programming. However, this method has two shortcomings. First, directly solving FIGRET requires a predicted traffic matrix. However, due to the existence of highly bursty source-destination pairs, finding an appropriate traffic prediction is difficult. Second, linear programming involves high computational complexity and may not scale to large networks. To address these issues, FIGRET leverages a deep neural network to accelerate TE computation. Similar to DOTE (Perry et al., 2023), FIGRET directly maps a history of traffic patterns to a routing weight configuration, thus eliminating the need of traffic prediction. To handle the path sensitivity constraints, FIGRET adds an additional term in its loss function to capture the customized robustness requirement.
@@
We conduct a comprehensive evaluation of FIGRET. This evaluation utilizes publicly available WAN datasets (Uhlig et al., 2006), as well as Data Center PoD-level and ToR-level topologies and traffic data (Alizadeh et al., 2013; Roy et al., 2015). The data encompasses topologies ranging from dozens to hundreds of nodes, with the corresponding traffic data exhibiting various characteristics, including profiles with low, moderate, and high burstiness. Through our evaluation, we find that FIGRET consistently delivers high-quality TE solutions across a variety of topologies. Compared to the TE system in Google’s production data center (Poutievski et al., 2022), FIGRET achieves an average Max Link Utilization (MLU) reduction of 9%-34% across different topologies and improves solution speed by 35×35$\times$35 ×-1800×1800$\times$1800 ×. In comparison with the state-of-the-art Deep Learning-based TE system DOTE (Perry et al., 2023), specifically designed for MLU optimization, FIGRET achieves notable improvements in two topologies with bursty traffic data. It reduced the average Max Link MLU by 4.5% and 5.3% while simultaneously decreasing the number of significant congestion incidents caused by traffic bursts by 41% and 53.9%. Meanwhile, in topologies with stable traffic data, FIGRET performs at least as well as DOTE, despite of the additional consideration of robustness. Finally, we numerically interpret FIGRET’s superior performance. Our code is available at (Fig, 2024).
@@
2.1. Necessity of managing bursts
Traffic engineering has been adopted in both Wide Area Networks (e.g., Google’s B4 (Jain et al., 2013) and Microsoft’s SWAN (Hong et al., 2013)) and data centers (e.g., Google’s optical data center network (Poutievski et al., 2022)) to improve network utilization and prevent network congestion.
To motivate the need to manage bursts in TE, we present an analysis of the impact of bursts on network performance and summarize the results in Figure LABEL:fig:_anti-burst_motivation. We implement two strategies in this study: 1) the ‘No hedging’ strategy, which uses the current traffic matrix to determine the TE configuration for the next interval without any approach to manage bursts; 2) the ‘Hedging’ strategy, which uses the current traffic matrix for configuring the next interval but incorporates the anti-burst Hedging mechanism utilized by Google in their Jupiter data center network (Poutievski et al., 2022). The underlying principle of the Hedging mechanism is to spread a flow across multiple paths to prevent bursts from excessively impacting any single path. We conduct evaluations on the GEANT WAN (Uhlig et al., 2006), and the PoD/ToR-level direct connect topologies [^1] We change the topology from clos topologies to direct-connect topologies, because TE is rarely used in clos topologies. As optical circuit switching witnessed widespread adoption in Google’s data centers (Poutievski et al., 2022), direct-connect topologies have become popular. with traffic traces collected in Meta’s data centers (Roy et al., 2015). Since GEANT WAN only offers minute-level traffic matrices, for fair comparison, we also aggregate Meta’s data center traffic trace at minute-level intervals. Our findings from Figure LABEL:fig:_anti-burst_motivation include:
  • Performance sensitivity to network variability: From the GEANT WAN network to the PoD-level data center network, and then to the ToR-level data center network, traffic becomes more volatile, and the performance of the ‘No hedging’ strategy becomes increasingly unstable.
  • Necessity of burst resistance: Across WAN and data center networks at both the ToR and PoD levels, the ‘No hedging’ strategy results in higher peaks on the MLU curve, indicating that the networks experience congestion due to bursts if anti-burst strategies are not employed.
  • Performance trade-offs with burst resistance: The ‘No hedging’ strategy displays higher peaks and lower troughs on the MLU curve. This can be interpreted as peaks occurring during bursts and troughs occurring in their absence. Conversely, the ‘Hedging’ strategy does not achieve as low troughs during non-burst conditions, because it forces a significant portion of traffic to take non-optimal paths.
In summary, managing bursts is necessary for TE, yet strategies for burst management often compromise non-burst scenario performance. So we seek a TE method that effectively manages bursts while minimizing the impact on non-burst scenario performance.
@@ 2.2. Diversity in traffic characteristics
We find that although traffic bursts occur, the extent of these bursts varies among different source-destination pairs (SD pairs). To demonstrate the diverse traffic characteristics under different SD pairs, we provide an analysis of traffic characteristics across SD pairs in various production networks. This includes the WAN network GEANT (Uhlig et al., 2006), as well as the Meta’s data center network (Roy et al., 2015) at both the PoD-level and ToR-level. Details on the public data used are provided in §5. For each SD pair, we calculate the variance of the traffic volumes as a measure of the dynamic changes in traffic. The results of our analysis are displayed in Figure LABEL:fig:example_of_traffic_demand, which shows the variance of traffic demand for different SD pairs across the three types of network traffic. The greater the variance for a SD pair, the more unstable that SD pair is.
The results in Figure LABEL:fig:example_of_traffic_demand indicate that, regardless of the topology type, the traffic characteristics are distinct for different SD pairs. If one treats the traffic from all the SD pairs equally, the resulting TE solution may either result in suboptimal performance in non-bursty scenarios or sacrifice the burst-handling capability. Hence, effectively leveraging the diverse traffic characteristics in TE is crucial for better balancing the performance trade-off between bursty and non-bursty situations.
@@ 2.3. Dig deep into the trade-off
TE scheme 1 considers all traffic as non-bursty and only optimizes for the normal situation. To minimize congestion in non-bursty scenarios, TE scheme 1 directs all traffic along the shortest paths. In the normal situation, its Max Link Utilization (MLU) is $\max\{\frac{1}{2},\frac{1}{2},\frac{1}{2}\}=0.5$. When any burst situation occurs, the MLU is increased to $\max\{\frac{4}{2},\frac{1}{2},\frac{1}{2}\}=2$.
TE scheme 2 considers all traffic as bursty traffic and aims to enhance robustness in response to the three burst situations by splitting traffic across different paths. When dealing with the normal situation, the MLU of TE scheme 2 is $\max\{\frac{1\times 0.5+1\times 0.5+1\times 0.5}{2},\frac{1\times 0.5+1\times 0.5+1\times 0.5}{2},\frac{1\times 0.5+1\times 0.5+1\times 0.5}{2}\}=0.75$, while dealing with the three burst situations, its MLU is $\max\{\frac{4\times 0.5+1\times 0.5+1\times 0.5}{2},\frac{4\times 0.5+1\times 0.5+1\times 0.5}{2},\frac{4\times 0.5+1\times 0.5+1\times 0.5}{2}\}=1.5$. That is, TE scheme 2 exhibits better robustness in handling traffic bursts compared to TE scheme 1, but with a decrease in normal performance.
Overall, this one-size-fits-all approach either leads to a network lacking the capacity to handle bursts or enhances the capability to manage bursts but at the cost of compromised performance in non-burst situations.
@@
Based on the observations in §2.1 and the insights from Figure LABEL:fig:example_of_traffic_demand and Figure LABEL:fig:TE_simple, it is evident that enhancing robustness to manage unexpected traffic bursts is necessary. However, uniformly increasing the robustness across all paths may lead to a compromise in average performance. A more nuanced strategy would be to adjust the robustness of paths based on the traffic characteristics of the traffic demands they serve. Compared to previous TE methods, this fine-grained robustness enhancement strategy in TE enables a better balance between normal-case performance and burst-case performance.
@@ 3. TE model
  • Network paths. Each source vertex $s$ communicates with each destination vertex $d$ via a set of network paths $P_{sd}$. We can also say that $P_{sd}$ serves the SD pair $(s,d)$.
  • Path capacity. The capacity of a path $p$ is denoted by $C_{p}$. It is the minimum capacity across the edges along the path (Poutievski et al., 2022).
  • TE configuration. A TE configuration $\mathcal{R}$ specifies for each source vertex $s$ and destination vertex $d$ how the $D_{sd}$ traffic from $s$ to $d$ is split across the paths in $P_{sd}$. A TE configuration specifies for each path $p\in P_{sd}$ a split ratio $r_{p}$, where $r_{p}$ is the fraction of the traffic demand from $s$ to $d$ forwarded along path $p$. The split ratios must satisfy $\sum_{p\in P_{sd}}r_{p}=1$.
@@
    Given a demand matrix $D$ and TE configuration $\mathcal{R}$, the total traffic traversing an edge $e$ is $f_{e}=\sum_{s,d\in V,p\in P_{sd},e\in p}D_{sd}\cdot r_{p}$, and then the MLU induced by $D$ and $\mathcal{R}$ is $\text{MLU}=\max_{e\in E}\frac{f_{e}}{c(e)}$, which we denote as $M(\mathcal{R},D)$.
TE under traffic uncertainty. At epoch $t$, the decision maker in TE must output a network configuration $\mathcal{R}_{t}(D_{1},\dots,D_{t-1})$ based on historical data $\{D_{1},\dots,D_{t-1}\}$ before the arrival of $D_{t}$. The objective is to obtain a network configuration that yields low link over-utilization for the upcoming traffic $D_{t}$.
@@
When configuring $\mathcal{R}_{t}$ based on historical data $\{D_{1},\dots,D_{t-1}\}$ before the arrival of $D_{t}$, it is critical to consider both anticipated and unanticipated traffic. An anticipated traffic demand $D_{t}^{\text{expect}}$ can be deduced from historical traffic data using certain traffic prediction technologies. However, we caution that such a prediction may not be accurate due to the existence of highly bursty traffic demands. To attain better network performance, it is crucial to incorporate the mismatch between the predicted traffic demand $D_{t}^{\text{expect}}$ and the real traffic demand $D_{t}$ when calculating each TE configuration.